OpenStudy (anonymous):

Sketch the graph of the equation (x+5)2 + y2 = 16 how would you graph this?

OpenStudy (anonymous):

solve for y first. y= mx+b

OpenStudy (anonymous):

its clear its open shape ,its ends goes to infinity best way .. set x=0 and get y ( 0,y) and set y =0 and get x (x,0) then take random points between ... x= 1,2,3,4,5,6, etc y = and plot good luck

OpenStudy (precal):

This is the equation of a circle center is (-5,0) and radius 2

OpenStudy (anonymous):

Precal .. its not square .. cause x not to the power 2. the bracts multiplied by 2

OpenStudy (precal):

I have a strong feeling the asker did not input the equation in correctly. I believe it is \[\left( x+5 \right)^{2}+y^2=16\]